# Deep Copula-Based Survival Analysis for Dependent Censoring with Identifiability Guarantees

### v3

@@ 1 Introduction
One recent approach to account for the dependencies is to assume that their true dependencies are known to the practitioners (Emura et al. 2018; Deresa and Van Keilegom 2021; Gharari et al. 2023). These methods utilize copulas, which are powerful statistical tools which model dependencies between random variables in isolation from their marginals. That is, if the copula between observed and censored times is known, then survival marginals may be unbiasedly estimated even in the presence of censored data. These methods face two significant and interrelated challenges. Firstly, specifying an underlying copula is inherently difficult, as practitioners may lack experience or prior information to make an informed choice. Secondly, misspecified copulas will exacerbate model bias, leading to incorrect inference and misleading results. More recent methods by Deresa and Van Keilegom alleviate these issues by learning the association parameters of the copula. Nonetheless, the problem of misspecifying the copula’s parametric form still exists.
@@ 4 Model Identifiability
(Czado and Van Keilegom 2022) Suppose the following conditions are satisfied: (C1) For all $\theta_{T_{1}},\theta_{T_{2}}\in\Theta_{T}$ and $\theta_{U_{1}},\theta_{U_{2}}\in\Theta_{U}$, we have
@@
(Czado and Van Keilegom 2022) Condition (C1) is satisfied by the families of Weibull, log-normal, log-logistic, log-Student, and Cox proportional hazard densities.
Theorem 3 and 4 are attributed to (Czado and Van Keilegom 2022). As it turns out, Archimedean copulas represented by $\varphi_{nn}$ have the attractive property of satisfying the condition (C2).
@@ 6 Related Work
Discussion on dependent censoring in survival analysis can be traced back to the seminal work of Tsiatis (1975) and Lagakos (1979). To account for dependent censoring, several copula-based approaches have been explored by assuming the family of ground truth copula is known. Czado and Van Keilegom (2022); Deresa and Van Keilegom proposed statistical methods for parametric survival marginals with identifiability guarantee. Midtfjord, Bin, and Huseby (2022) and Gharari et al. (2023) proposed to use boosting and neural network for maximizing the log-likelihood with pre-specified copulas, respectively. To the best of our knowledge, all existing copula-based approaches require practitioners to provide the ground truth copula, or at least its parametric form. Our proposed DCSurvival algorithm relaxes this assumption by approximating the copula with neural networks and learning their parameters.
@@ Appendix E Theorem Proofs
Theorem 4.
Suppose that the following two conditions are satisfied:
(C1) For $\theta_{T_{1}},\theta_{T_{2}}\in\Theta_{T}$ and $\theta_{U_{1}},\theta_{U_{2}}\in\Theta_{U}$, we have that
$\lim\limits_{t\to 0}\frac{f_{U,\theta_{U_{1}}}(t)}{f_{U,\theta_{U_{2}}}(t)}=1\Longleftrightarrow\theta_{U_{1}}=\theta_{U_{2}}.$
(C2) The parameter space $\Theta_{T}\times\Theta_{U}\times\Theta$ satisfies that
$\lim_{t\to 0}\left.\frac{\partial}{\partial u_{1}}\mathcal{C}_{\theta}(u_{1},u_{2})\right|_{\begin{subarray}{c}u_{1}=S_{T,\theta_{T}}(t)\\ u_{2}=S_{U,\theta_{U}}(t)\end{subarray}}=1$ (19)
$\lim_{t\to 0}\left.\frac{\partial}{\partial u_{2}}\mathcal{C}_{\theta}(u_{1},u_{2})\right|_{\begin{subarray}{c}u_{2}=S_{T,\theta_{T}}(t)\\ u_{2}=S_{U,\theta_{U}}(t)\end{subarray}}=1$ (20)
Then, the copula survival model in the main text is identified.
Proof.
The proof consists of two main parts. In the first part we show that the marginal distributions of the event and censoring times are identified by tying the densities of observed event/censoring times to the true density of event and censoring distributions. In a sense, this has similarity to showing that the observed outcomes equal to the potential outcomes by expectation in causal inference. However, unlike causal inference where the equivalence is achieved through confounding adjustment; in survival analysis we utilize the limit properties of the copulas. In the second part we show that the copula parameters are identified utilizing the uniqueness of copulas.
We now discuss identifiability of the survival marginals. From the distribution function of samples that have observed events we have
[…]
we can write the density of the observed samples that has experienced event as
$f_{y,\delta=1}(t)$ $=f_{T}(t)\cdot\Pr(U>t|T=t)$
$=f_{T}(t)\cdot S_{U|T}(t|t)$
$=f_{T}(t)\cdot\left.\frac{\partial}{\partial u_{1}}\mathcal{C}(u_{1},u_{2})\right|_{\begin{subarray}{c}u_{1}=S_{T}(t)\\ u_{2}=S_{U}(t)\end{subarray}}$ (22)
From condition (C2) we know that $\lim_{t\to 0}\partial\mathcal{C}(u_{1},u_{2})/\partial u_{1}=1$ when $u_{1}=S_{T}(t)$ and $u_{2}=S_{U}(t)$. Therefore, the densities of observed samples with $\delta=1$ equals to the ground truth densities of the event time when $t\to 0$, i.e.,
$\lim_{t\to 0}f_{y,\delta=1}(t)=\lim_{t\to 0}f_{T}(t),$ (23)
Suppose there are two sets of parameters $\theta_{T_{1}}$ and $\theta_{T_{2}}$ such that $f_{y,\delta=1,\theta_{T_{1}}}=f_{y,\delta=1,\theta_{T_{2}}}$ for all $t$, then
$1=\lim_{t\to 0}\frac{f_{y,\delta=1,\theta_{T_{1}}}(t)}{f_{y,\delta=1,\theta_{T_{2}}}(t)}=\lim_{t\to 0}\frac{f_{T,\theta_{T1}}(t)}{f_{T,\theta_{T2}}(t)}$ (24)
Given Condition (C1), we have
$\lim\limits_{t\to 0}\frac{f_{T,\theta_{T_{1}}}(t)}{f_{T,\theta_{T_{2}}}(t)}=1\Longleftrightarrow\theta_{T_{1}}=\theta_{T_{2}}.$ (25)
It follows that $\theta_{T1}=\theta_{T2}$ when $f_{y,\delta=1,\theta_{T1}}(t)=f_{y,\delta=1,\theta_{T2}}(t)$ for all $t$. Therefore, the identification of the marginal distributions of the event times is proved. The identification of the marginal distributions of the censoring times, i.e., $\theta_{U1}=\theta_{U2}$ when $f_{Y,\delta=0,\theta_{U1}}=f_{Y,\delta=0,\theta_{U2}}$ for all $t$, can be similarly shown.
We now demonstrate that the copula parameters are identified by showing that $\theta_{1}=\theta_{2}$ when $F_{y,\theta_{1}}(y)=F_{y,\theta_{2}}(y)$. Observing that for the observed time $Y=\min\{T,U\}$, we have
$F_{Y}(y)$ $=1-\Pr(Y>y)$
$=1-\Pr(T>y,U>y)$
$=1-\mathcal{C}(S_{T},S_{U}).$ (26)
Suppose there are two sets of parameters $\theta_{1}$ and $\theta_{2}$ such that $F_{y,\theta_{1}}(y)=F_{y,\theta_{2}}(y)$ for all $y$, then it follows that $1-C_{\theta_{1}}(S_{T},S_{U})=1-C_{\theta_{2}}(S_{T},S_{U})$. Noting that Sklar’s theorem states that the copula $\mathcal{C}$ is unique, we have $\theta_{1}=\theta_{2}$. Therefore, the parameters of the underlying copula are identified. ∎

### v4

@@ 1 Introduction
One recent approach to account for the dependencies is to assume the parametric form of the dependencies and learn the correlation parameter (Emura et al. 2018; Deresa and Van Keilegom 2021; Gharari et al. 2023). These methods utilize copulas, which are powerful statistical tools which model dependencies between random variables in isolation from their marginals. That is, if the copula between observed and censored times is known, then survival marginals may be unbiasedly estimated even in the presence of censored data. These methods face two significant and interrelated challenges. Firstly, specifying the parametric family of the underlying copula is inherently difficult, as practitioners may lack experience or prior information to make an informed choice. Secondly, misspecified copulas will exacerbate model bias, leading to incorrect inference and misleading results. More recent methods by Deresa and Van Keilegom alleviate these issues by learning the association parameters of the copula. Nonetheless, the problem of misspecifying the copula’s parametric form still exists.
@@ 4 Model Identifiability
(Czado and Van Keilegom 2023) Suppose the following conditions are satisfied: (C1) For all $\theta_{T_{1}},\theta_{T_{2}}\in\Theta_{T}$ and $\theta_{U_{1}},\theta_{U_{2}}\in\Theta_{U}$, we have
@@
(Czado and Van Keilegom 2023) Condition (C1) is satisfied by the families of Weibull, log-normal, log-logistic, and log-Student densities.
Theorem 3 and 4 are attributed to (Czado and Van Keilegom 2023). As it turns out, Archimedean copulas represented by $\varphi_{nn}$ have the attractive property of satisfying the condition (C2).
@@ 6 Related Work
Discussion on dependent censoring in survival analysis can be traced back to the seminal work of Tsiatis (1975) and Lagakos (1979). To account for dependent censoring, several copula-based approaches have been explored by assuming the parametric family of the ground truth copula is known. Specifically, Czado and Van Keilegom (2023); Deresa and Van Keilegom (2021, 2024) proposed statistical methods for parametric and semiparametric survival marginals with identifiability guarantee for all parameters. Midtfjord, Bin, and Huseby (2022) and Gharari et al. (2023) proposed to use boosting and neural network for maximizing the log-likelihood with pre-specified copulas, respectively. To the best of our knowledge, all existing copula-based approaches require practitioners to provide the ground truth copula, or at least its parametric form. Our proposed DCSurvival algorithm relaxes this assumption by approximating the copula with neural networks and learning their parameters.